OpenStudy (anonymous):

Find the vertices and foci of the ellipse given by the equation 4x^2+9y^2=144. a.)vertices: (0,6),(0,-6); foci: (0,2\sqrt5),(0,-2\sqrt5) b.)vertices: (4,0),(-4,0); foci: (2\sqrt5,0),(-2\sqrt5,0) c.)vertices: (0,4),(0,-4); foci: (0,2\sqrt5),(0,-2\sqrt5) d.) vertices: (6,0),(-6,0); foci: (2\sqrt5,0),(-2\sqrt5,0)

OpenStudy (mertsj):

Divide both sides by 144 to get it into the right form.

OpenStudy (anonymous):

im getting decimals.

OpenStudy (mertsj):

\[\frac{4x^2}{144}=\frac{x^2}{36}\]

OpenStudy (anonymous):

what happened to the y?

OpenStudy (mertsj):

I was only trying to show you what to do so you wouldn't get decimals. Now do what I said in the first place: divide both sides by 144